Find the length of major axis of an ellipse that has the distance between the foci equal to 10 and the length of the minor axis equal to 24.
Answer by ewatrrr(24785)

a^2 = sqrt(25 + 144) = 169
a = 13
the length of major axis of this ellipse is 26